OpenStudy (anonymous):

Intergrate. sqrt(5-4x-x^2)

OpenStudy (anonymous):

\[\int\limits_{}^{}\sqrt{5-4x-x^2}\]

OpenStudy (anonymous):

Looks like you should be able to complete the square, then use trig-sub

OpenStudy (anonymous):

\(\large \sqrt{5-4x-x^2}=\sqrt{-1[x^2+4x-5]}=\sqrt{-1[(x^2+4x+4)-4-5]} \) \(\large =\sqrt{-1[(x+2)^2-9]}=\sqrt{9-(x+2)^2}=\sqrt{3^2-(x+2)^2} \)

OpenStudy (anonymous):

Thank you so much!